# reference.bib

### v1

@@ 5. Manipulation of the Eisenstein Series
It is now clear from the above, that the analytic properties of the Dirichlet series of interest reduce to the ones of the Klingen Eisenstein series, as defined in Definition 4.1. This section is devoted in writing this Eisenstein series in a form of an Epstein zeta function, similar to [krieg1991, (7)]. In our case, because of the form of the Eisenstein series, we cannot use the method of Krieg in [krieg1991] with the minors of determinant. It turns out we can write the Eisenstein series in such a form, provided that the number of one-dimensional cusps is 11 11.
@@
$\mathcal{C}^{1}(\Gamma_{S})=\{\Gamma_{S}W\mid W\textup{ is isotropic in }V_{1}\}.$
A two-dimensional isotropic plane, or isotropic plane, is defined by two linearly independent vectors $g,h$. We normalise $g,h$ such that $g,h\in L_{1}^{*}$ and such that $\gcd{(S_{1}g)}=\gcd{(S_{1}h)}=1$. The isotropy condition means $S_{1}\left[\begin{pmatrix}l&m\end{pmatrix}\right]=0$.
@@
=$\mathfrak{H}=\{R\in M_{n+4}(\mathbb{R}):R=R^{t},RS_{1}^{-1}R=S_{1},\,\,R>0\}.$
The main proposition of the section then is the following:
@@
Let $S$ be such that #$\#\mathcal{C}^{1}(\Gamma_{S})=1$. Then, for each $Z\in\mathcal{H}_{S}$, there is $R_{Z}\in\mathfrak{H}$ such that
@@
Here, a matrix being primitive means that its elementary divisors are all 11 11 (see [Shimura_Euler_Product, Section 3]).
The rest of the section is devoted in proving this proposition. We start with two lemmas regarding the elements of $G_{\mathbb{R}}$, i.e. the orthogonal group attached to $S_{1}$.
@@
then by the relation $S_{1}[\gamma]=S_{1}$, we have
@@
Let $\gamma\in G_{\mathbb{R}}$ and write $\gamma=\begin{pmatrix}*\\ l^{t}\\ m^{t}\end{pmatrix}=\begin{pmatrix}*&*&*&*&*\\ l_{1}&l_{2}&x^{t}&l_{n+3}&l_{n+4}\\ m_{1}&m_{2}&y^{t}&m_{n+3}&m_{n+4}\end{pmatrix}$. Then if
@@
We now want to choose a specific majorant for $S_{1}$, so that the terms of the Eisenstein series take the form in Proposition 5.1. We start with defining the majorant for the identity element $I:=(i,0\cdots,0,i)^{t}\in\mathcal{H}_{S}$.
@@
Let $I:=(i,0,\cdots,0,i)^{t}$ and $R_{I}=\textup{diag}(1,1,S,1,1)$. Then, if $\gamma\in G_{\mathbb{R}}$, with $\gamma=\begin{pmatrix}*&l&m\end{pmatrix}^{t}$, we have
$\left(\frac{\textup{Im}(\gamma I)_{2}}{Q_{0}[\textup{Im}(\gamma I)]}\right)^{2}=\det{R_{I}[\ell]},$
@@
$\textup{Im}((\gamma I)_{2})=\textup{Im}\left(\frac{l_{1}+(l_{2}+l_{n+3})i+l_{n+4}}{m_{1}+(m_{2}+m_{n+3})i+m_{n+4}}\right)=\frac{(l_{2}+l_{n+3})(m_{1}+m_{n+4})-(l_{1}+l_{n+4})(m_{2}+m_{n+3})}{(m_{1}+m_{n+4})^{2}+(m_{2}+m_{n+3})^{2}}.$
@@
$Q_{0}[\textup{Im}(\gamma I)]=|j(\gamma,I)|^{-2}Q_{0}[\textup{Im}I]=\frac{2}{(m_{1}+m_{n+4})^{2}+(m_{2}+m_{n+3})^{2}}.$
@@
$\frac{\textup{Im}(\gamma I)_{2}}{Q_{0}[\textup{Im}(\gamma I)]}=(l_{2}+l_{n+3})(m_{1}+m_{n+4})-(l_{1}+l_{n+4})(m_{2}+m_{n+3}).$
On the other hand, we have
$R_{I}[\ell]=\begin{pmatrix}m_{1}^{2}+m_{2}^{2}+S^{-1}[y]+m_{n+3}^{2}+m_{n+4}^{2}&m_{1}l_{1}+m_{2}l_{2}+y^{t}S^{-1}x+m_{n+3}l_{n+3}+m_{n+4}l_{n+4}\\ l_{1}m_{1}+l_{2}m_{2}+x^{t}S^{-1}y+l_{n+3}m_{n+3}+l_{n+4}m_{n+4}&l_{1}^{2}+l_{2}^{2}+S^{-1}[x]+l_{n+3}^{2}+l_{n+4}^{2}\end{pmatrix}.$
But by the conditions we have, we get
$\begin{cases}S^{-1}[x]=2(l_{n+4}l_{1}+l_{n+3}l_{2}),\\ S^{-1}[y]=2(m_{n+4}m_{1}+m_{n+3}m_{2}),\\ x^{t}S^{-1}y=l_{n+4}m_{1}+l_{n+3}m_{2}+l_{2}m_{n+3}+l_{1}m_{n+4},\end{cases}$
so after taking determinant the claim follows. ∎
After defining the majorant for $I$, we can use the transitivity of the action defined in Section 2, in order to define a majorant for every $Z$ in $\mathcal{H}_{S}$.
@@
Let $Z\in\mathcal{H}_{S}$ and $\gamma=\begin{pmatrix}*&l&m\end{pmatrix}^{t}\in G_{\mathbb{R}}$. Then, $\exists R_{Z}\in\mathfrak{H}$ such that for all $\gamma\in G_{\mathbb{R}}$,
$\left(\frac{\textup{Im}(\gamma Z)_{2}}{Q_{0}[\textup{Im}(\gamma Z)]}\right)^{2}=\det(R_{Z}[\ell]),$
@@
We start the proof by constructing such a $R_{Z}$. We denote by $I=(i,0,\cdots,0,i)^{t}$. By transitivity, $\exists\delta\in G_{\mathbb{R}}^{0}$ such that $\delta\langle I\rangle=Z$. We then define $R_{Z}:=R_{I}[\delta^{-1}]$ and we claim this is well-defined. We prove this in Lemma 5.4. Then
$\left(\frac{\textup{Im}(\gamma Z)_{2}}{Q_{0}[\textup{Im}(\gamma Z)]}\right)^{2}=\left(\frac{\textup{Im}(\gamma\delta I)_{2}}{Q_{0}[\textup{Im}(\gamma\delta I)]}\right)^{2}=\det\left(R_{I}[\ell]\right),$
@@
$\left(\frac{\textup{Im}(\gamma Z)_{2}}{Q_{0}[\textup{Im}(\gamma Z)]}\right)^{2}=\det(R_{Z}[\delta\ell]).$
Now, if $\gamma=\begin{pmatrix}*\\ (l^{\prime})^{t}\\ (m^{\prime})^{t}\end{pmatrix}$, we want to show that the above quotient equals $\det(R_{Z}[S_{1}^{-1}\begin{pmatrix}m^{\prime}&l^{\prime}\end{pmatrix}])$. But
@@
Suffices to then show that (we remind here that $\ell=S_{1}^{-1}\begin{pmatrix}m&l\end{pmatrix}$)
@@
Finally, we give the final lemma, showing that $R_{Z}$ is well-defined.
@@
$R_{Z}$, as defined above, is well-defined.
@@
Using now equations $(2),(4),(6),(8)$, we can obtain
@@
Finally, from $(1),(6),(8)$ we have the equations
@@
  • Injective: If $\gamma^{-1}M=\delta^{-1}MN$ for some $N\in\textup{GL}_{2}(\mathbb{Z})$ then
@@
    which shows that $\gamma\delta^{-1}\in\Gamma_{S,J}$ (the first two columns of $\gamma\delta^{-1}$ belong in the $\mathbb{Z}$-span of the vectors $e_{1},e_{2}$ and $\gamma,\delta\in\Gamma_{S}\subset\hbox{SL}_{n+4}(\mathbb{Z})$).
  • Surjective: Consider now $\begin{pmatrix}l&m\end{pmatrix}\in X/\textup{GL}_{2}(\mathbb{Z})$. The plane generated by the vectors $l,m$ can be checked to be isotropic. Indeed, $\textup{span}\{l,m\}=\textup{span}\left\{\frac{l}{\gcd{(S_{1}l)}},\frac{m}{\gcd{(S_{1}m)}}\right\}$ in $V_{1}$. These vectors both belong in $L_{1}^{*}$ (can be checked directly) and also $\gcd{\left(\frac{S_{1}l}{\gcd{(S_{1}l)}}\right)}=1$ and similarly for the other vector. The isotropy condition still holds for the two new vectors (as we only change them by a scalar) and therefore the claim follows.
    Now, because #$\#\mathcal{C}^{1}(\Gamma_{S})=1$ by assumption, if $U$ is the plane generated by the basis vectors $e_{1},e_{2}$ ($U$ is an isotropic plane too), there is an element $K\in\Gamma_{S}$ such that $K$ maps $U$ to $W$. Hence, there exist $x,y,z,w\in\mathbb{Q}$ such that
@@
    Hence, as the inverse of that matrix also has integer entries, we must have that its determinant is ±$\pm 1$, i.e. $xw-yz=\pm 1$. Without loss of generality, we can assume the determinant is 11 11 (since we only work modulo $\textup{GL}_{2}(\mathbb{Z})$). We then consider the matrix
    $M_{D}=\begin{pmatrix}D^{*}\\ &1_{n}\\ &&D\end{pmatrix},$
    with $D=\begin{pmatrix}w&z\\ y&x\end{pmatrix}$, as after Definition 3.6. This belongs to $\Gamma_{S}$ (see [eisenstein_thesis, Example 1.3.4]). We can then check
    $KM_{D}e_{1}=K\begin{pmatrix}w\\ -y\\ \vdots\\ 0\end{pmatrix}=K(we_{1}-ye_{2})=\cdots=l,$
    and similarly $KM_{D}e_{2}=m$, i.e. the matrix $KM_{D}\in\Gamma_{S}$ has its first two columns equal to $l,m$ respectively. Now, $\gamma=(KM_{D})^{-1}\in\Gamma_{S}$ gets mapped to $\begin{pmatrix}l&m\end{pmatrix}$, as wanted.
The rest of the proof now follows by Proposition 5.2. ∎
Remark 1.
We would like to make a few comments here regarding the condition #$\#\mathcal{C}^{1}(\Gamma_{S})=1$. This condition is discussed in [eisenstein_thesis, Section 1.6.5]. However, the whole discussion there and the examples given (basically the Euclidean lattices, a complete list of the 7070 7070 that exist is known, see [nebe]) is about when $\Gamma_{S}$ is the integral connected component of the identity of $\hbox{O}(2,n+2)$ and not of $\hbox{SO}(2,n+2)$, as in our case. However, in the case when $\hbox{O}(S;\mathbb{Z})$ contains an element $K$ of determinant $-1$, we can show that there is a single orbit under the action of the SO-group too. Indeed, assume there is a single orbit under the action of the O-group. This means, that given an isotropic plane $W$ in $V_{1}$, $\exists\gamma\in\hbox{O}^{+}(L_{2})$ such that $\gamma W=\begin{pmatrix}e_{1}&e_{2}\end{pmatrix}$. If now $\gamma\in\hbox{SO}^{+}(L_{2})$, we are done. If not, consider the element
$\delta:=\begin{pmatrix}1\\ &1\\ &&K\\ &&&1\\ &&&&1\end{pmatrix}\in\hbox{O}^{+}(L_{2})$
of determinant $-1$. Then $\delta\gamma W=\delta\begin{pmatrix}e_{1}&e_{2}\end{pmatrix}=\begin{pmatrix}e_{1}&e_{2}\end{pmatrix}$ and $\gamma\delta\in\hbox{SO}^{+}(L_{2})$.
However, the question of whether such a $K$ exists is non-trivial, at least when $n$ is even (if $n$ is odd, $-1_{n}$ always works). A condition we can impose so that we ensure the existence of such an element can be found in [shimura2004arithmetic, Lemma 9.23, (iii)]. This is the existence of an element $x\in\mathbb{Z}^{n}$ so that $S[x]=2$. In our examples in practice, this is usually satisfied.
Finally, we have the following Lemma where we replace the condition of primitivity of the elements of $X$ in Proposition 5.1 with the condition that the elements have maximal rank.
@@
The proof is analogous to [deitmar_krieg, Lemma 3.1]. In particular, every matrix $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$ with rank =$\textup{rank }\ell=2$ and $S_{1}[\ell]=0$ can be written as $\ell=N\cdot M$ with $N$ being primitive and $S_{1}[N]=0$ and $M\in\textup{GL}_{2}(\mathbb{Q})\cap\textup{Mat}_{2,2}(\mathbb{Z})$. The proof then follows by the fact that
@@ 6. Theta Series and Transformation Properties
In order now to prove the analytic properties of the Klingen Eisenstein series $E(Z,s)$, we want to prove a theta correspondence between SO and $\hbox{Sp}_{2}$. That is, to integrate a Siegel Eisenstein series for $\hbox{Sp}_{2}$ against an appropriately defined theta series and end up with $E(Z,s)$. However, as in most cases, the inner-product integrals will diverge, so we need to apply an appropriate differential operator first. In this section, we recall the action of the symplectic group on Siegel’s upper half plane, define the appropriate theta series and prove some important transformation properties.
@@
$\hbox{Sp}_{m}(\mathbb{R})=\left\{g\in\hbox{GL}_{2m}(\mathbb{R})\mid g^{t}\begin{pmatrix}0_{m}&-1_{m}\\ 1_{m}&0_{m}\end{pmatrix}g=\begin{pmatrix}0_{m}&-1_{m}\\ 1_{m}&0_{m}\end{pmatrix}\right\}.$
@@
$\mathbb{H}_{m}=\left\{Z=X+iY\in M_{m}(\mathbb{C})\mid X=X^{t},Y=Y^{t}>0\right\}.$
Now $g=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\hbox{Sp}_{m}(\mathbb{R})$ acts on $\mathbb{H}_{m}$ via
@@
This defines a transitive action of $\hbox{Sp}_{m}(\mathbb{R})$ on $\mathbb{H}_{m}$. We also define the factor of automoprhy $j(\gamma,Z):=\det(CZ+D)$. We call $\hbox{Sp}_{m}(\mathbb{Z}):=\hbox{Sp}_{m}(\mathbb{R})\cap\hbox{GL}_{2m}(\mathbb{Z})$ the full modular group. For any $N>0$, we define the following congruence subgroup of $\hbox{Sp}_{m}(\mathbb{Z})$:
@@
Let $k\in\mathbb{Z}$. A function $F:\mathbb{H}_{m}\longrightarrow\mathbb{C}$, with $F\in\mathcal{C}^{\infty}(\mathbb{H}_{m})$, is called a $\mathcal{C}^{\infty}$-modular form of weight $k$ on the group $\Gamma_{0}^{(m)}(N)$ with a Dirichlet character $\psi\pmod{N}$, if for all
@@
$F(\gamma\langle Z\rangle)=\psi(\det D)\det(CZ+D)^{k}F(Z)$
@@
$\left\langle F,G\right\rangle:=\frac{1}{[\textup{Sp}_{m}(\mathbb{Z}):\Gamma_{0}^{(m)}(N)]}\int_{\Gamma_{0}^{(m)}(q)\backslash\mathbb{H}_{m}}F(Z)\overline{G(Z)}(\det Y)^{k}\hbox{d}^{*}Z,$
whenever the integral converges. Here, $\textup{d}^{*}Z$ is the $\hbox{Sp}_{m}(\mathbb{R})$-invariant measure
@@
We are now ready to give the definition of the theta series of interest, which is related to $\hbox{Sp}_{2}$. Such theta series were first studied by Siegel.
@@
Let also $\Theta(Z,W):=\det(Y)^{\frac{n+2}{2}}\theta(Z,W)$. Because $R_{W}$ is positive definite, it follows that $\Theta$ converges absolutely and uniformly in any region of the form $\mathbb{H}_{2}(\epsilon)=\{Z=X+iY\mid Y\geq\epsilon 1_{2}\}$ with $\epsilon>0$ and therefore defines a real analytic function of the variables $X,Y$ with $X+iY\in\mathbb{H}_{2}$ (cf. [spherical_theta, page 291]).
@@
The situation, however, is quite different when it comes to the transformation of $\Theta$ with respect to $\hbox{Sp}_{2}(\mathbb{Z})$. In particular, it is necessary to consider a specific congruence subgroup of $\hbox{Sp}_{2}(\mathbb{Z})$ and even then $\Theta$ transforms as a modular form of a non-trivial weight and character.
@@
Let $q$ denote the level of $S$, equivalently $S_{1}$ (see Definition 2.4). Then, for all $\gamma\in\Gamma_{0}^{(2)}(q)$, we have
@@
where $\chi_{S}(\gamma)$ is an eight root of unity which does not depend on $Z$ and $W$ (cf. [spherical_theta, Theorem 1]).

### v2

@@ 5. Manipulation of the Eisenstein Series
It is now clear from the above that the analytic properties of the Dirichlet series of interest reduce to the ones of the Klingen Eisenstein series, as given in Definition 4.2. This Section is devoted to writing this Eisenstein series in the form of an Epstein zeta function, similar to [krieg1991, equation (7)]. In our case, because of the form of the Eisenstein series, we cannot use the method of Krieg in [krieg1991] with the minors of the determinant. It turns out we can write the Eisenstein series in such a form, provided that the number of one-dimensional cusps is $1$.
@@
$\mathcal{C}^{1}(\Gamma_{S}):=\left\{\Gamma_{S}W\mid W\textup{ is a two-dimensional isotropic plane in }V_{1}\right\}.$
A two-dimensional isotropic plane, or isotropic plane, is defined by two linearly independent vectors $g,h\in V_{1}$. We normalise $g,h$ such that $g,h\in L_{1}^{*}$ and such that $\gcd{(S_{1}g)}=\gcd{(S_{1}h)}=1$. The isotropy condition means $S_{1}\left[\begin{pmatrix}g&h\end{pmatrix}\right]=0$.
@@
$\mathfrak{H}:=\{R\in M_{n+4}(\mathbb{R})\mid R=R^{t}>0,\ RS_{1}^{-1}R=S_{1}\}.$
The main Proposition of the Section is the following:
@@
Let $S$ be such that $\#\mathcal{C}^{1}(\Gamma_{S})=1$. Then, for each $Z\in\mathcal{H}_{S}$, there is $R_{Z}\in\mathfrak{H}$ such that
@@
Here, a matrix being primitive means that its elementary divisors are all $1$ (see [Shimura_Euler_Product, Section 3]).
The rest of the Section is devoted to proving this Proposition. We start with two Lemmas regarding the elements of $G_{\mathbb{R}}$, i.e., the special orthogonal group attached to $S_{1}$.
@@
then from the relation $S_{1}[\gamma]=S_{1}$, we have
@@
Let $\gamma\in G_{\mathbb{R}}$ and write $\gamma=\begin{pmatrix}*\\ l^{t}\\ m^{t}\end{pmatrix}=\begin{pmatrix}*&*&*&*&*\\ l_{1}&l_{2}&x^{t}&l_{n+3}&l_{n+4}\\ m_{1}&m_{2}&y^{t}&m_{n+3}&m_{n+4}\end{pmatrix}$. Then, if
@@
Now, for any $Z\in\mathcal{H}_{S}$, we want to choose a specific majorant for $S_{1}$, so that the terms of the Eisenstein series take the form in Proposition 5.3. We start with defining the majorant for the element $I:=(i,0\cdots,0,i)^{t}\in\mathcal{H}_{S}$.
@@
Let $I:=(i,0,\cdots,0,i)^{t}$ and $R_{I}:=\textup{diag}(1,1,S,1,1)$. Then, if $\gamma\in G_{\mathbb{R}}$, with $\gamma=\begin{pmatrix}*&l&m\end{pmatrix}^{t}$, we have
$\left(\frac{\textup{Im}((\gamma I)_{2})}{Q_{0}[\textup{Im}(\gamma I)]}\right)^{2}=\det{(R_{I}[\ell])},$
@@
$\textup{Im}((\gamma I)_{2})=\textup{Im}\left(\frac{l_{1}+(l_{2}+l_{n+3})i+l_{n+4}}{m_{1}+(m_{2}+m_{n+3})i+m_{n+4}}\right)=\\ =\frac{(l_{2}+l_{n+3})(m_{1}+m_{n+4})-(l_{1}+l_{n+4})(m_{2}+m_{n+3})}{(m_{1}+m_{n+4})^{2}+(m_{2}+m_{n+3})^{2}}.$
@@
$Q_{0}[\textup{Im}(\gamma I)]=|j(\gamma,I)|^{-2}Q_{0}[\textup{Im}I]=\frac{1}{(m_{1}+m_{n+4})^{2}+(m_{2}+m_{n+3})^{2}}.$
@@
$\frac{\textup{Im}((\gamma I)_{2})}{Q_{0}[\textup{Im}(\gamma I)]}=(l_{2}+l_{n+3})(m_{1}+m_{n+4})-(l_{1}+l_{n+4})(m_{2}+m_{n+3}).$
Now, from Lemma 5.5, with $\ell$ written as there, we have
$\begin{cases}S^{-1}[x]=2(l_{n+4}l_{1}+l_{n+3}l_{2}).\\ S^{-1}[y]=2(m_{n+4}m_{1}+m_{n+3}m_{2}).\\ x^{t}S^{-1}y=l_{n+4}m_{1}+l_{n+3}m_{2}+l_{2}m_{n+3}+l_{1}m_{n+4}.\end{cases}$
By then computing $R_{I}[\ell]$ and taking the determinant, the result follows.∎
After defining the majorant for $I$, we can use the transitivity of the action defined in (2), in order to define a majorant for every $Z$ in $\mathcal{H}_{S}$.
@@
Let $Z\in\mathcal{H}_{S}$. Then, $\exists R_{Z}\in\mathfrak{H}$ such that for all $\gamma\in G_{\mathbb{R}}$,
$\left(\frac{\textup{Im}((\gamma Z)_{2})}{Q_{0}[\textup{Im}(\gamma Z)]}\right)^{2}=\det(R_{Z}[\ell]),$
@@
We start the proof by constructing such an $R_{Z}$. We denote by $I=(i,0,\cdots,0,i)^{t}$. By transitivity, $\exists\delta\in G_{\mathbb{R}}^{0}$ such that $\delta\langle I\rangle=Z$. We then define $R_{Z}:=R_{I}[\delta^{-1}]$ and we claim this is well-defined. We prove this in Lemma 5.8. Then
$\left(\frac{\textup{Im}((\gamma Z)_{2})}{Q_{0}[\textup{Im}(\gamma Z)]}\right)^{2}=\left(\frac{\textup{Im}((\gamma\delta I)_{2})}{Q_{0}[\textup{Im}(\gamma\delta I)]}\right)^{2}=\det\left(R_{I}[\ell]\right),$
@@
$\left(\frac{\textup{Im}((\gamma Z)_{2})}{Q_{0}[\textup{Im}(\gamma Z)]}\right)^{2}=\det(R_{Z}[\delta\ell]).$
[…]
@@
[…]
@@
We now give the final Lemma, showing that $R_{Z}$ is well-defined.
@@
For each $Z\in\mathcal{H}_{S}$, $R_{Z}$, as defined in the proof of Proposition 5.7, is well-defined.
@@
Using now equations (4), (6), (8), (10), we obtain
@@
Finally, from (3), (8), (10), we have the equations
@@
  • Injective: If $\gamma^{-1}M=\delta^{-1}MN$ for some $N\in\textup{GL}_{2}(\mathbb{Z})$, then
@@
    which shows that $\gamma\delta^{-1}\in\Gamma_{S,J}$ (the first two columns of $\gamma\delta^{-1}$ belong in the $\mathbb{Z}$-span of the vectors $e_{1},e_{2}$ and $\gamma,\delta\in\Gamma_{S}\subset\textup{SL}_{n+4}(\mathbb{Z})$).
  • Surjective: Consider now $\begin{pmatrix}l&m\end{pmatrix}\in X/\textup{GL}_{2}(\mathbb{Z})$. The plane generated by the vectors $l,m$ can be checked to be isotropic. Indeed,
    $\textup{span}\{l,m\}=\textup{span}\left\{\frac{l}{\gcd{(S_{1}l)}},\frac{m}{\gcd{(S_{1}m)}}\right\}$
    in $V_{1}$. These vectors both belong in $L_{1}^{*}$ (can be checked directly) and also $\gcd{\left(\frac{S_{1}l}{\gcd{(S_{1}l)}}\right)}=1$ and similarly for the other vector. The isotropy condition still holds for the two new vectors (as we only change them by a scalar), and therefore the claim follows.
    Now, because $\#\mathcal{C}^{1}(\Gamma_{S})=1$ by assumption, if $U$ is the plane generated by the basis vectors $e_{1},e_{2}$ ($U$ is an isotropic plane too), there is an element $K\in\Gamma_{S}$ such that $K$ maps $U$ to $W$. Hence, there exist $x,y,z,w\in\mathbb{Q}$ such that
@@
    Hence, as the inverse of that matrix also has integer entries, we must have that its determinant is $\pm 1$, i.e., $xw-yz=\pm 1$. Therefore, $K^{-1}$ gets mapped to $\begin{pmatrix}l&m\end{pmatrix}\textup{GL}_{2}(\mathbb{Z})$, as wanted.
The rest of the proof now follows from Proposition 5.7. ∎
Remark 5.9.
We would like to make a few comments here regarding the condition $\#\mathcal{C}^{1}(\Gamma_{S})=1$. From [one-dimensional-cusps, Theorem 5.10], and because the lattice $L_{1}$ is maximal, we have that the condition $\#\mathcal{C}^{1}(\Gamma_{S})=1$ is equivalent to $L$ (the definite lattice of rank $n$) having only one class in its genus, and $[\textup{O}(L):\textup{SO}(L)]=2$. For $n\geq 2$, a complete list of lattices with one class in their genus is known (see [single-class-genera], [hanke2011enumerating] for example), and their rank is at most $10$. If $n=1$, $L=\mathbb{Z}$ also has a single class in its genus in our setting ($S=2t$, with $t\geq 1$ a square-free integer). Finally, for the condition $[\textup{O}(L):\textup{SO}(L)]=2$, a criterion we can use so that we ensure the existence of an element with determinant $-1$ (at least when $n$ is even as if $n$ is odd, $-1_{n}$ always works) can be found in [shimura2004arithmetic, Lemma 9.23, (iii)]. That is, the existence of some $x\in\mathbb{Z}^{n}$ so that $S[x]=2$.
Finally, we have the following Lemma, where we replace the condition of primitivity of the elements of $X$ in Proposition 5.3 with the condition that the elements have maximal rank.
@@
The proof is analogous to [deitmar_krieg, Lemma 3.1]. In particular, every matrix $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$ with $\textup{rank }\ell=2$ and $S_{1}[\ell]=0$ can be written as $\ell=N\cdot M$ with $N$ being primitive and $S_{1}[N]=0$ and $M\in\textup{GL}_{2}(\mathbb{Q})\cap\textup{Mat}_{2,2}(\mathbb{Z})$. The proof then follows from the fact that
@@ 6. Theta Series and Transformation Properties
In order now to prove the analytic properties of the Klingen Eisenstein series $E(Z,s)$, we want to prove a theta correspondence between $\textup{SO}(2,n+2)$ and $\textup{Sp}_{2}$. That is, to integrate a Siegel Eisenstein series for $\textup{Sp}_{2}$ against an appropriately defined theta series and end up with $E(Z,s)$. However, as in most cases, the inner-product integrals will diverge, so we need to apply an appropriate differential operator first. In this Section, we recall the action of the symplectic group on Siegel’s upper half plane, define the appropriate theta series and prove some important transformation properties.
@@
$\textup{Sp}_{m}(\mathbb{R}):=\left\{g\in\textup{GL}_{2m}(\mathbb{R})\mid g^{t}\begin{pmatrix}0_{m}&-1_{m}\\ 1_{m}&0_{m}\end{pmatrix}g=\begin{pmatrix}0_{m}&-1_{m}\\ 1_{m}&0_{m}\end{pmatrix}\right\}$.
@@
$\mathbb{H}_{m}:=\left\{Z=X+iY\in M_{m}(\mathbb{C})\mid X=X^{t},Y=Y^{t}>0\right\}.$
Now $g=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\textup{Sp}_{m}(\mathbb{R})$ acts on $\mathbb{H}_{m}$ via
@@
This defines a transitive action of $\textup{Sp}_{m}(\mathbb{R})$ on $\mathbb{H}_{m}$. We also define the factor of automorphy $j(g,Z):=\det(CZ+D)$. We call $\textup{Sp}_{m}(\mathbb{Z}):=\textup{Sp}_{m}(\mathbb{R})\cap\textup{GL}_{2m}(\mathbb{Z})$ the full modular group. For any integer $N>0$, we define the following congruence subgroup of $\textup{Sp}_{m}(\mathbb{Z})$:
@@
Let $k\in\mathbb{Z}$ and $m\geq 2$. A function $F:\mathbb{H}_{m}\longrightarrow\mathbb{C}$, with $F\in\mathcal{C}^{\infty}(\mathbb{H}_{m})$, is called a $\mathcal{C}^{\infty}$-modular form of weight $k$ on the group $\Gamma_{0}^{(m)}(N)$ with a Dirichlet character $\psi\pmod{N}$, if for all
@@
$F(\gamma\langle Z\rangle)=\psi(\det D)\det(CZ+D)^{k}F(Z),$
@@
$\left\langle F,G\right\rangle:=\frac{1}{[\textup{Sp}_{m}(\mathbb{Z}):\Gamma_{0}^{(m)}(N)]}\int_{\Gamma_{0}^{(m)}(N)\backslash\mathbb{H}_{m}}F(Z)\overline{G(Z)}(\det Y)^{k}\textup{d}^{*}Z,$
whenever the integral converges. Here, $\textup{d}^{*}Z$ is the $\textup{Sp}_{m}(\mathbb{R})$-invariant measure
@@
We are now ready to give the definition of the theta series of interest, which is related to $\textup{Sp}_{2}$. Such theta series were first studied by Siegel.
@@
Let also $\Theta(Z,W):=\det(Y)^{\frac{n+2}{2}}\theta(Z,W)$. Because $R_{W}$ is positive definite, it follows that $\Theta$ converges absolutely and uniformly in any region of the form $\mathbb{H}_{2}(\epsilon)=\{Z=X+iY\mid Y\geq\epsilon 1_{2}\}$ with $\epsilon>0$ and therefore defines a real analytic function of the matrices $X,Y$ with $X+iY\in\mathbb{H}_{2}$ (cf. [spherical_theta, page 291]).
@@
The situation, however, is quite different when it comes to the transformation of $\Theta$ with respect to $\textup{Sp}_{2}(\mathbb{Z})$. In particular, it is necessary to consider a specific congruence subgroup of $\textup{Sp}_{2}(\mathbb{Z})$ and even then $\Theta$ transforms as a modular form of a non-trivial weight and character.
@@
Let $q$ denote the level of $S$, equivalently $S_{1}$ (see Definition 2.4). Then, for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma_{0}^{(2)}(q)$, we have
@@
where $\chi_{S}(\gamma)$ is an eighth root of unity which does not depend on $Z$ and $W$ (cf. [spherical_theta, Theorem 1]).